A body of a 6kg undergoes a Constant horizontal acceleration of 2 m² 2. Calculate the resultant horizontal force acting on the body. what will be the resultant force on the body when it moves with a uniform velocity of 5m/s?.


1
Expert's answer
2021-10-18T10:59:00-0400

Given:

"m=6\:\rm kg"

"a_1=2\:\rm m/s^2"

"a_2=0\:\rm m/s^2"


The Newton's second law says

"F=ma"

Hence,

(1)

"F=6*2=12\:\rm N"

(2)

"F=6*0=0\:\rm N"